3.3. If in a ring R every x ∈ R satisfies x2 = x, prove that R

must be commutative

(A ring in which x2 = x for all elements is called a Boolean ring).


1
Expert's answer
2022-06-16T09:40:52-0400

Let "x,y\in R". Then "(x + y)^2 = (x + y)(x + y) = x^2 + xy + yx + y^2"

Since "x^2 = x" and "y^2 = y" we have "x + y = x + xy + yx + y".

Hence "xy = \u2212yx".

But for every "x\in R": "(\u2212x) = (\u2212x)^2 = (\u2212x)(\u2212x) = x^2 = x".

Hence "\u2212yx = yx" i.e. we obtain "xy = yx".